# The Value of Nothing111“Nowadays people know the price of everything and the value of nothing,” The Picture of Dorian Gray Wilde (1890): Multimodal Extraction of Human Values Expressed by TikTok Influencers

### v1

@@ Abstract
Societal and personal values are transmitted to younger generations through interaction and exposure. Traditionally, children and adolescents learned values from parents, educators, or peers. Nowadays, social platforms serve as a significant channel through which youth (and adults) consume information, as the main medium of entertainment, and possibly the medium through which they learn different values. In this paper we extract implicit values from TikTok movies uploaded by online influencers targeting children and adolescents. We curated a dataset of hundreds of TikTok movies and annotated them according to the Schwartz Theory of Personal Values. We then experimented with an array of Masked and Large language model, exploring how values can be detected. Specifically, we considered two pipelines – direct extraction of values from video and a 2-step approach in which videos are first converted to elaborated scripts and then values are extracted.
Achieving state-of-the-art results, we find that the 2-step approach performs significantly better than the direct approach and that using a trainable Masked Language Model as a second step significantly outperforms a few-shot application of a number of Large Language Models. We further discuss the impact of fine-tuning and compare the performance of the different models on identification of values present or contradicted in the TikTok. Finally, we share the first values-annotated dataset of TikTok videos. Our results pave the way to further research on influence and value transmission in video-based social platforms.
@@ 1 Introduction
The transmission of values between generations is a key process for the preservation of society (Schönpflug 2001). Classic research assumed that transmission of values is carried out mainly by parents (Grusec 2011) or schools (Berson and Oreg 2016). However, youth spend hours every day on social media (Smahel et al. 2020), where they encounter new role models (Tolbert and Drogos 2019). The social media environment has been shown to strongly influence children’s behavior, in the context of consumer behavior (Coates et al. 2019). However, influencers in social media promote more than products: they form communities, by establishing connections, encouraging engagement, and promoting shared values (Himelboim and Golan 2023).
@@
The Schwartz Theory of Personal Values (Schwartz 1992) offers a well-studied framework for the investigation of human motivation. It defines values as abstract goals, describing the end states aspired by people to achieve in their lives, such as success, independence, and care. These goals cover a wide range of motivations, with different individuals and societies, varying in the prioritization they place on different values (Schwartz 1992; Sagiv and Schwartz 2022). The theory identifies 19 core values, characterized by the goals they promote (Schwartz 2012). The values and their definitions, as used in the shared taskes at SemEval-2023 (Kiesel et al. 2022) and CLEF-2024 (Kiesel et al. 2024b) are provided in Table 1. Those 19 values are expressed through a wide variety of judgments, evaluations, and behaviors (Sagiv and Roccas 2021). Personal values, as defined by Schwartz, were consistently identified in different contexts, such as national cultures throughout the world (Sagiv and Schwartz 2022), organizational cultures (Daniel, Arieli, and Akerman 2024), traditional media (Kiesel et al. 2024a), and online disagreements (van der Meer et al. 2023).
@@
In this work we suggest that social media influencers targeting youth can be characterized not only by genre, topic, and style, but also by the values they pursue and promote. For example, influencers hosting competitions between followers promote values of achievement, those presenting a unique craft project promote creativity, and others who pull pranks on strangers promote disregard for social norms. A specific example, presented in Table 2, depicts a gaming influencer failing to return tennis balls served by a professional tennis player. The values manifested in the video are Achievement (trying to properly return fast serves) and conflicted Face (the influencer show himself failing again and again to no avail).
@@
How can we identify the values transmitted to youth through social media? This question has only recently began to be addressed. Some studies identified values prevalent in written social media using methods ranging from LIWC word matching to masked language models (e.g., BERT) (Kumar et al. 2018; Silva, Lo, and Lim 2020). However, recent trends lean heavily toward visual based social media, such as TikTok (Smahel et al. 2020; Rejeb et al. 2024). The shift toward visual mediums introduces further challenge, since it requires analyzing a clip, combining text, audio, and video. Further, the clips are typically very short, providing content through little stimuli. To the best of our knowledge, this is the first attempt to extract values from TikTok specifically, and visual social media in general.
In this work we experiment with an array of algorithms to extract values from TikTok videos targeting adolescents. To that end, we have curated a unique annotated corpus of 890 TikToks, posted by influencers targeting children and adolescents, covering a variety of genres prevalent in TikTok (see Section 3 for details). The corpus is useful for the investigation of values prevalent in this platform, and will be open for analysis.
@@
We further experiment with and compare methods of value extraction, including fine-tuned Masked Language Models and Large Language Models. We go on to evaluate each method against the manual annotations, to identify the most effective one. Specifically, we consider two main pipelines: a direct extraction from LLMs, leveraging their multimodal capabilities, and a 2-step extraction approach in which a video is first converted to a detailed textual script (master scene heading, action, dialogue) and the values are extracted from the text, rather than directly from the video. We achieve a higher F-score than previously reported state-of-the-art results (Schroter, Dementieva, and Groh 2023; Legkas et al. 2024), in spite the more challenging setting (a direct comparison is unattainable, see Section 2). We find that that the 2-step approach perform better than the direct approach and that using a trainable Masked Language Model in the second step significantly outperforms a few-shot application of a Large Language Model. We further discuss the impact of fine-tuning and compare the performance of the different models on the present manifestation of values and the conflicted manifestations. Finally, we share the first values-annotated dataset of TikTok videos.
@@ 2 Related Work
TikTok and Influencers Research
Youth spend hours on various social media sites (Smahel et al. 2020; Tome and Cardita 2021), with TikTok gaining increased popularity (Anderson, Faverio, and Gottfried 2023; Woodward 2024). TikTok is characterized by engaging short videos that can include pervasive use of filters, and background music (Rejeb et al. 2024). The use of a costume “for you” page and “like” button increases immersion in the application (Montag, Yang, and Elhai 2021).
Social media in general, and TikTok specifically, are characterized by a large number of Social Media Influencers (SMIs). Social media influencers are social media users who gained large following (Ki et al. 2020) by creating and curating content that portrays an appealing online personality (Tafesse and Wood 2021). Influencers form parasocial relationships with their followers: members of the audience feels close, as if they know the influencers intimately (Reinikainen et al. 2020). Although parasocial relationships are known within traditional mass media, they may increase in the context of social media, because of its interactive nature (Conde and Casais 2023; Lou and Yuan 2019), and the disclosure of the influencers’ personal lives (Conde and Casais 2023). Followers often identify with their favorite influencers (Lou and Yuan 2019), making them likely role models for young viewers.
Ample research supports the understanding that children and adolescents can learn from mass traditional media (Prot et al. 2015), as well as social media (Ragelienė and Grønhøj 2021; Craig, Brooks, and Bichard 2021). From the media, youth can acquire information (Dorey and McCool 2009), but also form their worldviews (Duffett 2017). A recent study established that the following of social media influencers is associated with higher importance ascribed to materialistic values among Finish youth (Tuominen et al. 2023). We suggest that in order to further investigate the role of influencers in youth value transmission, we must identify the values they embrace.
@@
Values are abstract goals describing end-states individuals aspire to achieve in their lives (Sagiv and Schwartz 2022; Schwartz 1992). Values direct a wide variety of behaviors, ranging between life-altering choices, such as career preferences (Lipshits-Braziler, Arieli, and Daniel 2024), and daily behaviors, such as the choice of touristic destinations (Saji, Vij, and Kabiraj 2015).
@@
The Schwartz Theory of Personal Values is the leading theory of values in social psychology, well-validated in hundreds of samples across the world (Sagiv and Roccas 2017). Importantly, the theory extends beyond moral values, such as justice and care, to define a variety of motivations that can motivate individuals (Twito-Weingarten and Knafo‐Noam 2022), including, for example, safety, excitement, and influence. Moreover, the values are organized in a dynamic system of inter-relations, as the pursuit of one value may contrast the pursuit of conflicting values, yet enhance the pursuit of compatible ones (Schwartz 1992). These inherent trade-offs require individuals to prioritize among values. The inter-relations can be summarized into two main axes. One describes the opposition between self-transcendence (care for others, close and distant) and self-enhancement (focus on advancement of self-interest). The second describes the opposition between conservation (preservation of the status-quo) and openness to change (embracing new experiences).
@@
Research established that children and adolescents hold meaningful values, following the content and structure defined by the Schwartz Theory of Personal Values (Daniel, Misgav, and Chomsky 2024; Daniel and Benish-Weisman 2019; Knafo-Noam, Daniel, and Benish-Weisman 2023). It is often agreed that children’s values are mainly learned from the social environment, in a process of value transmission (Twito and Knafo-Noam 2020). However, the social environment is varied and includes many socialization agents. Such agents are parents (Grusec 2011; Knafo-Noam, Barni, and Schwartz 2020), schools (Berson and Oreg 2016; Auer et al. 2023), the peer group (Benish-Weisman, Oreg, and Berson 2022; Benish‐Weisman 2024), and the national culture (Daniel et al. 2012). As social media becomes an important factor in the lives of adolescents, it is crucial to learn whether individuals within this sphere, such as influencers, become socialization agents teaching youth personal values.
Values in the Media
@@
Schwartz values were identified in content displayed across social media platforms, such as Instagram (Trillò et al. 2021), Reddit (Chen et al. 2014), Facebook and Twitter (Kumar et al. 2018). Values in social media content were presented in varied ways, such as memes (Shifman 2019), and new-years’ resolutions (Hallinan et al. 2023). However, little is known about the prevalence of Schwartz values in youth-directed media. Some research exists on the prevalence of values in child-directed mass traditional media. An analysis of television shows addressing 9-11 year olds, showed high representation of values of fame (Uhls and Greenfield 2011). Similarly, television shows addressing young children, displayed work as an expression of a large variety of values, including benevolence, security, conformity, achievement, universalism, stimulation and hedonism (Aharoni, Te’eni-Harari, and Eyal 2024). Our work aims at opening the door to expand this work by developing an AI-based value classification in TikTok.
@@
The field of automated value detection has evolved significantly. Focusing on methods for classifying values according to Schwartz’s Theory of Personal Values (Schwartz 2012), early approaches leveraged the Linguistic Inquiry and Word Count (LIWC) (Pennebaker et al. 2015), or a targeted dictionaries (Ponizovskiy et al. 2020) detecting values through lexical mathcing. Values were detected in texts produced across various platforms, including Twitter (Kadić-Maglajlić, Lages, and Pantano 2024), narrative texts (Fischer et al. 2022), song lyrics (Demetriou et al. 2024), and social media platforms such as Reddit (Oulahbib, Benhaddi, and El Hadaj 2024) and Facebook (Kumar et al. 2018). Notable advances include the development of stack models combining LIWC features trained on Facebook data with cross-platform validation on Twitter (Silva, Lo, and Lim 2020). VALUENET (Qiu et al. 2022), introduced a large-scale text dataset and implemented a Transformer-based value regression model for human value modeling in dialogue systems.
@@
Recent developments in predicting values were introduced with the application of masked langauge models. The SemEval-2023 (Kiesel et al. 2022) and CLEF-2024 (Kiesel et al. 2024b) competitions focused on value detection in argumentative texts and multi-lingual traditional media contexts respectively. These competitions demonstrated the effectiveness of BERT-based architectures (Schroter, Dementieva, and Groh 2023; Legkas et al. 2024) in capturing nuanced value expressions across two different text genres and languages. It also demonstrated that the algorithm developed in one context was not successful in identification of values in a second context (Kiesel et al. 2024b), asserting the importance of contextual models in value detection.
@@
Generative AI methods were hardly utilized for value detection, with the exception of the FULCRA project (Yao et al. 2024), leveraging LLM-generated text pairs for safety-oriented value detection. Our work extends these methodological advances by pioneering value detection in TikTok’s unique multimodal environment. Unlike previous approaches that focused on single modalities, TikTok content presents a complex interplay of written text, audio, and visual elements. To address this challenge, we conduct a comparative analysis of two leading approaches: BERT-based architectures and LLM-based methods. By evaluating these distinct methodologies on the same TikTok dataset, we provide the first systematic comparison of their effectiveness in detecting values within short-form video content. This novel application to multimodal social media content, combined with our comparative analysis, can contribute valuable insights to both social media analysis and automated value detection research, particularly for platforms where meaning is conveyed through multiple channels simultaneously.
Figure 1: Values per movie in the annotated data.
@@
Figure 2: Occurrences of each value in the annotated dataset. Blue bars indicate the appearance of a value (e.g., achievement) and red bars indicate the conflicted value (e.g., apathy and lack of ambition, conflicting with achievement).
@@ 3.1 Data Curation
We identified potential TikTok influencers using surveys of children and adolescents (anonymous survey, approved by a the local IRB), and lists provided by three large language models (ChatGPT, Gemini, Bing). Two research assistants reviewed the potential influencers, to confirm they are targeting children or adolescents and to classify them by genre. This process resulted in a list of 89 influencers, producing content in the major genres prevalent in child- and adolescent-targeted TikTok: beauty and skincare, lifestyle, entertainment and pranks, crafts and DYI, vlogging, gaming and dancing, singing and lip-syncing. We sampled 10 TikToks from the page of each influencer. We first identified the pinned TikToks, and supplemented the sample with the last TikToks available on March 19, 2024, until a total of 10 TikToks was reached from each influencer, to the total N = 890.
We only selected TikToks that include verbal sound, such as talking, narration or song. The crawling and curation were done with slight modifications to the API shared by (Teather 2024).
3.2 Annotation Protocol
We constructed a codebook on the basis of the Schwartz theory of personal values (Schwartz 1992, 2012). The codebook included a definition to each of the 19 values derived from the theory of personal values (Schwartz 2012). In addition, the theory included examples of expression of these values that are common on TikTok. The codebook included instructions for identifying values as present (1, marked as +), when it is expressed or pursued in the TikTok; as absent (0), when the TikTok does not refer to the value; or as conflicted (-1, marked as -), when the TikTok contradicts or hinders the pursuit of the value. Each TikTok could be annotated as referring to as many or as few values as relevant. A team of annotators was trained in value annotation. Each TikTok was independently annotated by two annotators. A third annotator consolidated the annotations by resolving discrepancies. To assess inter-rater agreement on our nominally scaled data, we considered various metrics beyond traditional $\pi$ and $\kappa$ statistics, which are known to be sensitive to imbalanced datasets (Gwet 2008). Given that our dataset exhibits significant imbalance, with the majority category (not present) accounting for 0.76 of annotations and rare categories (conflict and dominant) comprising only 0.06 and 0.03, we calculated Gwet’s AC1 coefficient, which has been demonstrated to be more robust against prevalence effects and marginal probability issues than Cohen’s Kappa (Wongpakaran et al. 2013). Our analysis yielded an AC1 score of 0.77, indicating substantial inter-rater agreement despite the challenging data distribution. The codebook is available at Appendix A.1.
In total, the annotated dataset of 890 TikTok TikToks, were annotated with a total of 2968 labels (of values and conflicting values). The distribution of values per movie is presented in Figure 1 and the frequency of each value in our dataset is presented in Figure 2. We make the annotated dataset available at [URL to added in camera ready version].

### v2

# The Value of Nothing111“Nowadays people know the price of everything and the value of nothing,” The Picture of Dorian Gray (Wilde 1890): Multimodal Extraction of Human Values Expressed by TikTok Influencers

@@ Abstract
Societal and personal values are transmitted to younger generations through interaction and exposure. Traditionally, children and adolescents learned values from parents, educators, or peers. Nowadays, social platforms serve as a significant channel through which youth (and adults) consume information, as the main medium of entertainment, and possibly the medium through which they learn different values. In this paper we extract implicit values from TikTok movies uploaded by online influencers targeting children and adolescents. We curated a dataset of hundreds of TikTok movies and annotated them according to the well established Schwartz Theory of Personal Values. We then experimented with an array of language models, investigating their utility in value identification. Specifically, we considered two pipelines: direct extraction of values from video and a 2-step approach in which videos are first converted to elaborated scripts and values are extracted from the textual scripts.
We find that the 2-step approach performs significantly better than the direct approach and that using a few-shot application of a Large Language Model in both stages outperformed the use of a fine-tuned Masked Language Model in the second stage. We further discuss the impact of continuous pretraining and fine-tuning and compare the performance of the different models on identification of values endorsed or confronted in the TikTok. Finally, we share the first values-annotated dataset of TikTok videos.
To the best of our knowledge, this is the first attempt to extract values from TikTok specifically, and visual social media in general. Our results pave the way to future research on value transmission in video-based social platforms.
@@ 1 Introduction
The transmission of values between generations is a key process for the preservation of society (Schönpflug 2001). Classic research assumed that transmission of values is carried out mainly by parents (Grusec 2011) or schools (Berson and Oreg 2016). However, youth spend hours every day on social media (Smahel et al. 2020), where they find new role models (Tolbert and Drogos 2019). The social media environment has been shown to strongly influence children’s behavior, in the context of consumer behavior (Coates et al. 2019). However, influencers on social media promote more than products: they form communities and promote shared values (Himelboim and Golan 2023).
@@
The Schwartz Theory of Personal Values (Schwartz 1992) offers a well-studied framework for the investigation of human motivation. It defines values as abstract goals, describing the end states people aspire to achieve in their lives, such as success, independence, and care. These goals cover a wide range of motivations, with different individuals and societies varying in the prioritization they place on different values (Schwartz 1992; Sagiv and Schwartz 2022). Values are expressed through a wide variety of judgments, evaluations, and behaviors (Sagiv and Roccas 2021). The theory identifies 19 core values, characterized by the boader goals they convey or serve (Schwartz 2012). Personal values, as defined by Schwartz (see Table 1), were consistently identified in different contexts, such as national cultures throughout the world (Sagiv and Schwartz 2022), organizational cultures (Daniel et al. 2024a), traditional media (Kiesel et al. 2024a), and online discourse (van der Meer et al. 2023).
@@
In this work, we suggest that social media influencers targeting youth, can be characterized not only by genre, topic, and style, but also by the values they endorse and the values they confront and disregard. To list a number of examples, an influencer hosting a competition between followers endorses values of achievement; an influencer presenting a unique craft project endorses values of creativity; influencers who pull pranks on strangers confront the values of respect for social norms. A number of values can be endorsed/confronted in the same piece of content. For example, in one video (script available in Table 2), a gaming influencer fails to return tennis balls served by a professional tennis player. The value of Achievement (trying to properly return fast serves) is endored in the video. At the same time, the value of Face is confronted, as the influencer is humiliated (shows himself failing again and again to no avail).
@@
How can we identify the values transmitted to youth through social media? Some recent studies identified values prevalent in written social media using a variety of methods e.g. (Kumar et al. 2018; Silva et al. 2020) ranging from linguistic inquiry and word count (LIWC), word matching to masked language models, such as BERT. However, current trends in social media lean heavily toward visual based media, such as TikTok (Smahel et al. 2020; Rejeb et al. 2024). The shift toward visual mediums introduces a further challenge, since it requires analyzing a clip in a multimodal way, combining text, audio, and video. In addition, the clips are typically very short, providing rich content through limited stimuli. To the best of our knowledge, this is the first attempt to extract values from TikTok specifically, and visual social media in general.
@@
In this work we have curated and annotated a unique corpus of 885 TikToks, posted by influencers targeting children and adolescents, covering a variety of genres prevalent in TikTok (see Section 3 for details). The corpus is useful for the investigation of values prevalent in this platform, and will be open for analysis. We experiment with and compare an array of algorithms of value classification, including fine-tuned Masked Language Models and Large Language Models. Specifically, we consider two main pipelines: a direct extraction from LLMs, leveraging their multimodal capabilities, and a 2-step extraction approach in which a video is first converted to a detailed textual script (master scene heading, action, dialogue) and the values are extracted from the text, rather than directly from the video. In spite of the multimodal complexity, we achieve a high F-score, higher than the state-of-the-art results reported on the task of extracting values from texts (Schroter et al. 2023; Legkas et al. 2024).
We find that the 2-step approach performs better than the direct approach and that using a few-shot application of a Large Language Model outperforms a trainable Masked Language Model in the second step. We further discuss the impact of fine-tuning and compare the performance of the different models on the endorsed and the confronted values.
@@ 2 Related Work
The TikTok Influencers Ecosystem
Youth spend hours on various social media sites (Smahel et al. 2020), with TikTok gaining increased popularity (Anderson et al. 2023; Woodward 2024). TikTok is characterized by engaging short videos that can include widespread use of filters, and background music (Rejeb et al. 2024). The use of a custom “For You” page and “Like” button increases immersion in the application (Montag et al. 2021).
Social media in general, and TikTok specifically, are characterized by a large number of Social Media Influencers (SMIs). Social media influencers are social media users who gained a large following (Ki et al. 2020) by creating and curating content that portrays an appealing online personality (Tafesse and Wood 2021). Influencers form parasocial relationships with their followers: members of the audience feel close to influencers, as if they know them intimately (Reinikainen et al. 2020). Influencers are especially likely to form such relationships because of social media’s interactive nature (Conde and Casais 2023; Lou and Yuan 2019), and the disclosure of the influencers’ personal lives (Conde and Casais 2023). Followers often identify with their favorite influencers (Lou and Yuan 2019), making them likely role models for young viewers.
Ample research supports the understanding that children and adolescents can learn from mass traditional media (Prot et al. 2015), as well as social media (Craig et al. 2021). From the media, youth can acquire information (Dorey and McCool 2009), but also form their worldviews (Duffett 2017). A recent study established that the following of social media influencers is associated with higher importance ascribed to materialistic values among Finish youth (Tuominen et al. 2023). We suggest that in order to further investigate the role of influencers in youth value transmission, we must identify the values they embrace.
@@
Values are abstract goals describing end-states individuals aspire to achieve in their lives (Sagiv and Schwartz 2022; Schwartz 1992). Values direct a wide variety of behaviors, ranging between life-altering choices such as career preferences (Lipshits-Braziler et al. 2024), and daily behaviors, such as the choice of touristic destinations (Saji et al. 2015).
@@
The Schwartz Theory of Personal Values is the leading theory of values in social psychology, well-validated in hundreds of samples across the world (Sagiv and Roccas 2017). Importantly, the theory extends beyond moral values, such as justice and care, to define a variety of motivations that can drive individuals (Twito-Weingarten and Knafo‐Noam 2022), including, for example, safety, excitement, and influence. Moreover, the values are organized in a dynamic system of inter-relations, as the pursuit of one value may contrast the pursuit of conflicting values, and enhance the pursuit of compatible ones (Schwartz 1992). These inherent trade-offs require individuals to prioritize among values. The inter-relations can be summarized into two main axes. One describes the tension between self-transcendence (care for others, close and distant) and self-enhancement (focus on advancement of self-interest). The second describes the tension between conservation (preservation of the status-quo) and openness to change (embracing new experiences). The list of Schwartz values and their short description is available at Table 1.
@@
Research established that children and adolescents hold meaningful values, following the content and structure defined by the Schwartz Theory of Personal Values (Daniel et al. 2024b; Daniel and Benish-Weisman 2019; Knafo-Noam et al. 2023). It is often agreed that children’s values are mainly learned from the social environment, in a process of value transmission (Twito and Knafo-Noam 2020). However, the social environment is varied and includes many socialization agents. Such agents are parents (Grusec 2011; Knafo-Noam et al. 2020), schools (Berson and Oreg 2016), the peer group (Benish-Weisman et al. 2022; Benish‐Weisman 2024), and the national culture (Daniel et al. 2012). As social media becomes an important factor in the lives of adolescents, it is crucial to learn whether individuals within this sphere, such as influencers, become socialization agents teaching personal values to youth.
Values Social Media
@@
Schwartz values were identified in content displayed across social media platforms, such as Instagram (Trillò et al. 2021), Reddit (Chen et al. 2014), Facebook and Twitter (Kumar et al. 2018). Values were presented in social media content in varied ways, such as memes (Shifman 2019), and new-years’ resolutions (Hallinan et al. 2023). However, little is known about the prevalence of Schwartz values in youth-directed media. Some research exists on the prevalence of values in child-directed mass traditional media. An analysis of television-shows addressing 9-11 year olds, showed high representation of fame values (Uhls and Greenfield 2011). Similarly, television shows addressing young children, displayed work as an expression of a large variety of values, including benevolence, security, conformity, achievement, universalism, stimulation and hedonism (Aharoni et al. 2024). Our work aims at opening the door to extract values at scale by developing an AI-based value detection in TikTok movies.
@@
The field of automated value detection has evolved significantly. Focusing on methods for classifying values according to Schwartz’s Theory of Personal Values (Schwartz 2012), early approaches leveraged the Linguistic Inquiry and Word Count (LIWC) (Pennebaker et al. 2015), or a targeted dictionary (Ponizovskiy et al. 2020) detecting values through lexical matching. Values were detected in texts produced across various platforms, including Twitter (Kadić-Maglajlić et al. 2024), narrative texts (Fischer et al. 2022), song lyrics (Demetriou et al. 2024), and social media platforms such as Reddit (Oulahbib et al. 2024) and Facebook (Kumar et al. 2018). Notable advances include the development of stack models combining LIWC features trained on Facebook data with cross-platform validation on Twitter (Silva et al. 2020). VALUENET (Qiu et al. 2022), introduced a large-scale text dataset and implemented a Transformer-based value regression model for human value modeling in dialogue systems.
@@
Recent developments in value prediction were introduced with the application of masked language models. The SemEval-2022 (Kiesel et al. 2022) and CLEF-2024 (Kiesel et al. 2024b) competitions focused on value detection in argumentative texts and multi-lingual traditional media contexts respectively. CLEF-2024 was the first to identify both endorsed and confronted values in texts. These competitions demonstrated the effectiveness of BERT-based architectures (Schroter et al. 2023; Legkas et al. 2024) in capturing nuanced value expressions across two different text genres and languages. It also demonstrated that algorithms developed in one context may not succeed in identification of values in a another context (Kiesel et al. 2024b), asserting the importance of contextual models in value detection.
@@
Generative AI methods have not yet been utilized for value detection, with the exception of the FULCRA project (Yao et al. 2024), leveraging LLM-generated text pairs for safety-oriented value detection. Our work extends these methodological advances by pioneering value detection in TikTok’s unique multimodal environment. Unlike previous approaches that focused on single modalities, TikTok content presents a complex interplay of written text, audio, and visual elements. To address this challenge, we conduct a comparative analysis of two leading approaches: BERT-based architectures and LLM-based methods. By evaluating these distinct methodologies on the same TikTok dataset, we provide the first systematic comparison of their effectiveness in detecting values within short-form video content. This novel application to multimodal social media content, combined with our comparative analysis, can contribute valuable insights to both social media analysis and automated value detection research, particularly for platforms where meaning is conveyed through multiple channels simultaneously.
Figure 1: Unique values per TikTok
@@
Figure 2: Occurrences of each value in the annotated dataset. Blue bars indicate the appearance of an endorsed value (e.g., achievement) and red bars indicate the confronted value (e.g., apathy and lack of ambition, conflicting with achievement).
@@ 3.1 Data Curation
We identified relevant TikTok influencers using anonymous surveys, asking children and adolescents (local IRB approved) to list the TikTok content creators they follow. We also prompted three language models (ChatGPT, Gemini, Bing) to provide lists of prominent influencers with clout over young users. Two RAs verified the accounts and confirmed they indeed target children or adolescentsand classified them by genre. This process resulted in a list of 91 influencers, active in the following genres: beauty and skincare, lifestyle, entertainment and pranks, crafts and DYI, vlogging, gaming and dancing, singing and lip-syncing. We sampled ten TikToks from the page of each influencer, including all TikToks pinned by the influencers, if existed. The full dataset contains $N=885$ movies.
The crawling and curation were done with slight modifications to the Teather API (Teather 2024), collecting only publicly available data. Such data curation, handling, processing and analysis, requires careful consideration of ethical aspects and potential harms, from privacy issues to intellectual property rights. We discuss potential harms, ethical considerations and handling protocols in Section 7.
3.2 Data Annotation
Annotation guidelines were based on the Schwartz theory of personal values (Schwartz 1992, 2012). The guidelines provide a definition for each of the 19 values, along with examples of how values are expressed. For each TikTok and value, annotators were asked to mark whether the values is endorsed (expressed or pursued in the TikTok; 1), absent (not referenced explicitly or implicitly; 0) or confronted (contradicted or hindered; -1). Each TikTok was independently annotated by two annotators from a team trained by the authors. A third annotator consolidated the annotations by resolving discrepancies. The annotation guidelines are available in Appendix A.1. We converted each TikTok annotation to a vector of 38 binary labels representing the 19 endorsed values and 19 confronted values.
The annotated dataset of 885 TikToks contains 2933 (endorsed/confronted) values. The distribution of values per TikTok is presented in Figure 1 and the frequency of each value in the dataset is presented in Figure 2. The annotated dataset (TikTok IDs and annotations) is available at [URL]. We used Gwet’s AC1 coefficient to assess inter-rater agreement. This coefficient estimates reliability beyond 2 rateres, and has been demonstrated robust against prevalence imbalance and marginal probability issues (Wongpakaran et al. 2013). The AC1 agreement was 0.77, indicating substantial inter-annotator agreement.